40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Aug 24 11:30:27 2011, jamie, jenne, kiwamu, suresh, steve, Update, SUS, ETMX wiped, replaced, door on 
    Reply  Thu Aug 25 08:12:09 2011, steve, Update, SUS, ETMY & ITMY wiped, replaced, vac door on 
       Reply  Thu Aug 25 13:10:42 2011, Jenne, Update, SUS, Drag wiping 
Message ID: 5295     Entry time: Wed Aug 24 11:30:27 2011     Reply to this: 5300
Author: jamie, jenne, kiwamu, suresh, steve 
Type: Update 
Category: SUS 
Subject: ETMX wiped, replaced, door on 

We've closed up ETMX:

  • the optic was drag wiped
  • the suspension tower was put back in place
  • earthquake stops were backed off the appropriate number of turns, and de-ionized
  • chamber door was put on
ELOG V3.1.3-